Main INDEX
Monthly INDEX
PREV
NEXT
User name George Lolos
Log entry time 15:53:04 on June17,2000
Entry number 44053
keyword=Day shift run summary
08:12......Run 2168 is finished..........Target cooling pressure is too high....Current is reduced.....Conditions inherited from previous shift. Unless noted otherwise in the log, present target condition will be labeled as normal.
08:12......Run 2169 started.
08:33...Beam OFF. 08:34...Beam ON. Beam spot (x&y) remains unchanged.
08:39......Run 2169 finished. Conditions were normal and beam was stable.
08:40......Run 2170 started.
09:07......Run 2170 ended.
09:07......Run 2171 started. Frantic phone exchanges with "cryo experts" and Adam have left us in the same state of understanding. Kuss is on his way.
09:33.....Run 2171 ended.
09:34......Run 2172 started. Kuss is working on the target. At 09:51 the Hadron Q1 tripped. Run terminated at 09:51.
09:56, Q1 was reset from upstairs. No end-station entry required.
09:57......Run 2173 started.
Hadron Q1 has reached nominal field value. Target operating but still a problem in the high pressure return line. Kuss is pursuing the matter further.
10:26.....Run 2173 ended.
10:26.....Run 2174 started.
10:53......Run 2174 ended.
10:53.....Run 2175 started.
Normal running, short beam interruptions. Kuss' evaluation of the transfer line problem is that we either have an erroneous reading or there is an obstraction. In the meantime, the target appears to behave normally and no restrictions have been placed on beam current. Asked MCC to give us between 70-80 uA.
11:24.....Run 2175 ended.
11:24.....Run 2176 started.
11:54.....Run 2176 ended.
11:55.....Run 2177 started. Beam intensity is now 70 uA.
12:23.....Run 2177 ended.
12:23.....Run 2178 started.
ESR temperature has decreased to 12K. Investigating with MCC.
12:52.....Run 2178 ended.
12:53.....Run 2179 started.
ESR temperature back to normal without human intervention. Another spooky element in the Hall A saga.
13:18.....Run 2179 ended.
13:19.....Run 2180 started.
ESR has been "fixed" and problem identified as a part of feedback loop when we adjusted the high pressure return line elements. Therefore, disregard the reference to spooky elements. Status is now normal.
13:47.....Run 2180 ended.
13:47...Run 2181 started.
14:13....Run 2181 ended.
14:14....Run 2182 started.
14:41....Run 2182 ended.
14:42....Run 2183 started.
15:12....Run 2183 ended.
15:12....Run 2184 started.
15:38....Run 2184 ended.
15:39....Run 2185 started.